Solve by substitution:3y+4x=5 x+y=1
Answer by rfer(16322)

3(1-x)+4x=5
3-3x+4x=5
x=2
------------
3y+4*2=5
3y+8=5
3y=-3
y=-1
-----------------
2-1=1